OpenStudy (anonymous):

if a card is drawn at random from a deck of 52 cards, what is the probability that the card (jack, queen, king) or a diamond?

OpenStudy (amistre64):

how many cards fit the bill?

OpenStudy (amistre64):

granted that the question assumes you know how a deck of playing cards is built to start with.

OpenStudy (elonasushchik):

So there are 4 suits...each has 13 cards in it....13 * 4 = 52 We want...either a black card....or a face card.....we know the face cards are jacks...queens...kings so there are 4 suits...and 3 face cards in each suit so 4 * 3 = 12 face cards... we also want a black card if not a face card....there are *not counting face cards now* 52 - 12 = 40 cards that are not face cards....there are only 2 suits that are red and 2 that are black...so we have 40 / 2 = 20 possible black cards... so altogether we have 20 black cards and 12 face cards 20 + 12 = 32 this is out of ALL cards so 3252 simplify this down to get...???

OpenStudy (elonasushchik):

Medal??

OpenStudy (amistre64):

hmm, your response is a good example but it should be noted that it does not use the original questions information.

OpenStudy (elonasushchik):

Yes. It's an exampl.